Carlas Pizza, a small pizzeria operating in a competitive market. Its short-run production function for the current plant isQuestion 5 Homework • Answered In the scenario above, if Carlas uses the expanded plant, what is the total cost of producingQuestion 6 Homework . Answered In the scenario above, suppose the going price of pizza is $8.35. What is the firms optimal l
Answer #1

Question 5. Total cost of 1099 pizza = ATC * Q = 9.38*1099 = 10308.62 or 10309

Question 6. We find MC from the table by using the total cost schedule. We see that when price is 8.35, MC is 8.35 at 1217 pizzas. Hence the correct choice is option E

Labor Quantity ATC in long run Total cost in long run Marginal cost
1 116 16.3 1890.8 16.3
2 290 11.97 3471.3 9.08
3 497 10.58 5258.26 8.63
4 717 9.93 7119.81 8.46
5 925 9.58 8861.5 8.37
6 1099 9.38 10308.62 8.32
7 1217 9.28 11293.76 8.35
8 1254 9.26 11612.04 8.60